# Single-photon scattering in giant-atom topological-waveguide-QED systems

### v1

@@ III.3.4 The ABBA-coupling configuration
For the ABBA-coupling configuration, the dimensionless parameters in Eqs. (35a)-(36f) are $\mu_{1(4)}=\nu_{2(3)}=1$ and $\nu_{1(4)}=\mu_{2(3)}=0$. In Figs. 8(a) and 8(b), we plot the reflection coefficient as a function of $\Delta_{k}$ and $\Delta$. Different from the previous three coupling cases, it can be seen that the reflection spectra have two reflection peaks ($R_{2}=1$) at $\Delta_{k}=\Gamma_{e}\sin(k_{\Delta}d-\phi_{k_{\Delta}})$ and $\Delta_{k}=\Gamma_{e}\sin(k_{\Delta}d+\phi_{k_{\Delta}})$. Here, the locations of the two peaks depend on the sign of $\delta$. To show more details, we plot in Figs. 8(c) and 8(d) the corresponding profiles of Figs. 8(a) and 8(b) at various values of $\Delta$, respectively.
To analyze the characteristics of the reflection spectra more clearly, we substitute $\delta=\pm 0.5$ and $\Delta\approx 1.20J$ (the red solid curves) into Eqs. (LABEL:r2)-(36f) and obtain the reflection rate
@@
Here, the characteristic parameters become $\Delta_{L1,k_{\Delta}}=\Delta_{L2,k_{\Delta}}=\Delta_{L,k_{\Delta}}\approx-0.96\Gamma_{e}$ ($0.46\Gamma_{e}$), $\Gamma_{1,k_{\Delta}}+\Gamma_{2,k_{\Delta}}=4\Gamma_{e}$, $g_{12,k_{\Delta}}\approx 0$, and $\Gamma_{1,k_{\Delta}}\Gamma_{2,k_{\Delta}}-\Gamma_{12,k_{\Delta}}^{2}\approx 0$ for $\delta=\pm 0.5$. Therefore, the reflection spectrum exhibits the Lorentzian line shapes centered at $\Delta_{k}=-0.96\Gamma_{e}$ ($0.46\Gamma_{e}$) with the linewidth of $4\Gamma_{e}$, as shown by the red solid curves in Figs. 8(c) and 8(d), respectively. When $\delta=0.5$ ($-0.5$) and $\Delta\approx 1.89J$, the characteristic quantities become $\Delta_{L1,k_{\Delta}}=\Delta_{L2,k_{\Delta}}=\Delta_{L,k_{\Delta}}\approx 0.98\Gamma_{e}$ ($0.83\Gamma_{e}$), $\Gamma_{1,k_{\Delta}}+\Gamma_{2,k_{\Delta}}=4\Gamma_{e}$, $g_{12,k_{\Delta}}\approx 0$, and $\Gamma_{1,k_{\Delta}}\Gamma_{2,k_{\Delta}}-\Gamma_{12,k_{\Delta}}^{2}\approx 0$, which satisfy Eq. (46), as shown by the purple dot-dashed curve in Fig. 8(c) [Fig. 8(d)]. In other parameters, the reflection spectra have two completely reflection peaks, such as $\Delta=1.40J$ (the blue dashed curves). In particular, when $\delta=0.5$ $(-0.5)$ and $\Delta\approx 1.58J$, we have $\Delta_{\text{S}L,k_{\Delta}}\approx\Delta_{\text{A}L,k_{\Delta}}\approx 0$, $\Gamma_{\text{S},k_{\Delta}}\approx 0$, $\Gamma_{\text{A},k_{\Delta}}\approx 7.79\Gamma_{e}$ $(5.26\Gamma_{e})$, $g_{\text{SA},k_{\Delta}}\approx-0.32\Gamma_{e}$ $(-0.95\Gamma_{e})$, and $\Gamma_{\text{SA},k_{\Delta}}\approx 0$. Here, the exchange interaction strength $g_{\text{SA},k_{\Delta}}$ and the individual decay rate of the antisymmetric state $\Gamma_{\text{A},k_{\Delta}}$ satisfy the condition $|g_{\text{SA},k_{\Delta}}|<\Gamma_{\text{A},k_{\Delta}}/4$ Feng2021 ; Abi-Salloum2010 ; Anisimov2011 . In this case, the system exhibits an electromagnetically induced transparency (EIT)-like spectrum. Thus, we find that the reflection spectrum has a minimal value ($R_{2}\approx 0$) at $\Delta_{k}\approx 0$, and that the reflection spectrum has two symmetric complete-reflection peaks ($R_{2}=1$) at $\Delta_{k}\approx\pm 0.32\Gamma_{e}$ (±$(\pm 0.95\Gamma_{e})$, as shown by the black dotted curve in Fig. 8(c) [Fig. 8(d)].
Figure 9: The reflection coefficient $R_{2}$ for the AAAB-coupling configuration as a function of $\Delta_{k}$ and $\Delta$. The dimerization parameter is (a) $\delta=0.5$ and (b) $\delta=-0.5$. The profiles of panels (a) and (b) at several values of $\Delta$ are shown by the curves in (c) and (d). Other parameters used are $g=0.01J$ and $d=2$.
@@ III.3.5 The AAAB-coupling configuration
For the AAAB-coupling configuration, the dimensionless parameters in Eqs. (35a)-(36f) are given by $\mu_{1-3}=\nu_{4}=1$ and $\nu_{1-3}=\mu_{4}=0$. In Figs. 9(a) and 9(b), we show the reflection coefficient as a function of $\Delta_{k}$ and $\Delta$. We find that the reflection spectra have two reflection peaks ($R_{2}=1$) at $\Delta_{k}=\Gamma_{e}\sin(k_{\Delta}d)$ (the first reflection peak) and $\Delta_{k}=\Gamma_{e}\sin(k_{\Delta}d-\phi_{k_{\Delta}})$ (the second reflection peak). This means that the position of the second reflection peak is influenced by the sign of $\delta$. To show the details clearly, in Figs. 9(c) and 9(d) we plot the profiles of Figs. 9(a) and 9(b) at different frequency detunings in the region $\Delta\in(J,2J)$, respectively.
When $\delta=\pm 0.5$ and $\Delta\approx 1.58J$ (the blue dashed curves), the reflection spectra exhibit the Lorentzian line shapes. Similarly, we substitute $\delta=\pm 0.5$ and $\Delta\approx 1.58J$ into Eqs. (LABEL:r2)-(36f) and obtain the reflection coefficient as
@@
Here, we have $\Delta_{L1,k_{\Delta}}=0$, $\Delta_{L2,k_{\Delta}}\approx-0.32\Gamma_{e}$ ($-0.95\Gamma_{e}$), $\Gamma_{1,k_{\Delta}}=0$, $\Gamma_{2,k_{\Delta}}\approx 3.9\Gamma_{e}$ ($2.6\Gamma_{e}$), $g_{12,k_{\Delta}}\approx 0$, and $\Gamma_{12,k_{\Delta}}\approx 0$ for $\delta=0.5$ ($-0.5$). Therefore, the first giant atom is decoupled from the waveguide, while the second giant atom is coupled to the waveguide. In this case, the reflection spectrum exhibits the Lorentzian line shape centered at $\Delta_{k}=-0.32\Gamma_{e}$ ($-0.95\Gamma_{e}$) with the linewidth of $3.9\Gamma_{e}$ ($2.6\Gamma_{e}$), as shown by the blue dashed curve in Fig. 9(c) [Fig. 9(d)]. In particular, when $\delta=\pm 0.5$, $\Delta\approx 1.12J$ (the red solid curves), $1.62J$ (the black dotted curves), and $1.70J$ (the purple dot-dashed curves), the number of the reflection peak depends on the sign of the dimerization parameter $\delta$. Specifically, when $\Delta\approx 1.12J$ (the red solid curves) and $1.70J$ (the purple dot-dashed curves), the reflection spectra at $\delta>0$ have two reflection peaks, while the reflection spectra at $\delta<0$ have one reflection peak, and vice versa at $\Delta\approx 1.62J$ (the black dotted curves). This is because the reflection peaks are located at $\Delta\approx 1.12J$ or $1.70J$ when $\delta=-0.5$ ($\Delta\approx 1.62J$ when $\delta=0.5$), i.e., $\Delta_{k}=\Gamma_{e}\sin(k_{\Delta}d)=\Gamma_{e}\sin(k_{\Delta}d-\phi_{k_{\Delta}})$. In this case, the reflection spectra have one reflection peak. These analyses indicate that only the position of the second peak depends on the sign of $\delta$. When the sign of $\delta$ changes, the position of the second peak changes accordingly.
@@ III.3.6 The AABA-coupling configuration
For the AABA-coupling configuration, the dimensionless parameters in Eqs. (35a)-(36f) are $\mu_{1,2,4}=\nu_{3}=1$ and $\nu_{1,2,4}=\mu_{3}=0$. Figures 10(a) and 10(b) show the reflection coefficient as a function of $\Delta_{k}$ and $\Delta$. It can be found that the reflection spectra have two peaks at $\Delta_{k}=\Gamma_{e}\sin(k_{\Delta}d)$ (the first reflection peak) and $\Delta_{k}=\Gamma_{e}\sin(k_{\Delta}d+\phi_{k_{\Delta}})$ (the second reflection peak), respectively. Therefore, only the position of the second peak depends on the sign of $\delta$, which is similar to the AAAB-coupling configuration. In Figs. 10(c) and 10(d) we show the profiles of Figs. 10(a) and 10(b) at different frequency detunings in the region $\Delta\in(J,2J)$, respectively.
Figure 10: The reflection coefficient $R_{2}$ for the AABA-coupling configuration as a function of $\Delta_{k}$ and $\Delta$. The dimerization parameter is (a) $\delta=0.5$ and (b) $\delta=-0.5$. The profiles of panels (a) and (b) at several values of $\Delta$ are shown by the curves in (c) and (d). Other parameters used are $g=0.01J$ and $d=2$.
@@
When $\delta=\pm 0.5$ and $\Delta\approx 1.58J$ (the black dotted curves), the reflection spectra exhibit the Lorentzian line shapes, which is consistent with the AAAB-coupling configuration. According to Eqs. (LABEL:r2)-(36f), we can obtain the reflection coefficient $R_{2}^{\text{AABA}}$, which satisfies Eq. (47). Here, the characteristic quantities become $\Delta_{L1,k_{\Delta}}=0$, $\Delta_{L2,k_{\Delta}}\approx 0.32\Gamma_{e}$ ($0.95\Gamma_{e}$), $\Gamma_{1,k_{\Delta}}=0$, $\Gamma_{2,k_{\Delta}}\approx 3.9\Gamma_{e}$ ($2.6\Gamma_{e}$), $g_{12,k_{\Delta}}\approx 0$, and $\Gamma_{12,k_{\Delta}}\approx 0$ for $\delta=0.5$ ($-0.5$). Therefore, the first giant atom is decoupled from the waveguide, while the second giant atom is coupled to the waveguide. In this case, the reflection spectrum exhibits the Lorentzian line shape centered at $\Delta_{k}=0.32\Gamma_{e}$ ($0.95\Gamma_{e}$) with the linewidth of $3.9\Gamma_{e}$ ($2.6\Gamma_{e}$), as shown by the black dotted curve in Fig. 10(c) [Fig. 10(d)]. From the above analyses, it can be seen that when $\delta=0.5$ ($-0.5$) and $\Delta\approx 1.58J$, the reflection spectra of AAAB- and AABA-coupling configuration are symmetric with respect to $\Delta_{k}=0$, as shown by the blue dashed curve in Fig. 9(c) [Fig. 9(d)] and the black dotted curve in Fig. 10(c) [Fig. 10(d)]. For other values of $\Delta$, such as $\Delta=1.70J$, the reflection spectra have two reflection peaks [see the purple dot-dashed curves in Figs. 10(c) and 10(d)]. In particular, when $\delta=\pm 0.5$, $\Delta\approx 1.38J$ (the red solid curves) and $1.54J$ (the blue dashed curves), the number of the reflection peaks depends on the sign of $\delta$. Specifically, when $\Delta\approx 1.38J$ (the red solid curves), the reflection spectrum at $\delta>0$ has two reflection peaks, while the reflection spectrum at $\delta<0$ has one reflection peak, and vice versa at $\Delta\approx 1.54J$ (the blue dashed curves). When $\delta=-0.5$ and $\Delta\approx 1.38J$ ($\delta=0.5$ and $\Delta\approx 1.54J$), we have the relation $\Delta_{k}=\Gamma_{e}\sin(k_{\Delta}d)=\Gamma_{e}\sin(k_{\Delta}d+\phi_{k_{\Delta}})$. Thus, the reflection spectra have one reflection peak, which is similar to the AAAB-coupling configuration.

### v2

@@ III.3.4 The ABBA-coupling configuration
For the ABBA-coupling configuration, the dimensionless parameters in Eqs. (35a)–(36f) are $\mu_{1(4)}=\nu_{2(3)}=1$ and $\nu_{1(4)}=\mu_{2(3)}=0$. In Figs. 8(a) and 8(b), we plot the reflection coefficient as a function of $\Delta_{k}$ and $\Delta$. Different from the previous three coupling cases, it can be seen that the reflection spectra have two reflection peaks ($R_{2}=1$) at $\Delta_{k}=\Gamma_{e}\sin(k_{\Delta}d-\phi_{k_{\Delta}})$ and $\Delta_{k}=\Gamma_{e}\sin(k_{\Delta}d+\phi_{k_{\Delta}})$. Here, the locations of the two peaks depend on the sign of $\delta$. To show more details, we plot in Figs. 8(c) and 8(d) the corresponding profiles of Figs. 8(a) and 8(b) at various values of $\Delta$, respectively.
To analyze the characteristics of the reflection spectra more clearly, we substitute $\delta=\pm 0.5$ and $\Delta\approx 1.20J$ (the red solid curves) into Eqs. (35b)–(36f) and obtain the reflection rate
@@
Here, the characteristic parameters become $\Delta_{L1,k_{\Delta}}=\Delta_{L2,k_{\Delta}}=\Delta_{L,k_{\Delta}}\approx-0.96\Gamma_{e}$ ($0.46\Gamma_{e}$), $\Gamma_{1,k_{\Delta}}+\Gamma_{2,k_{\Delta}}=4\Gamma_{e}$, $g_{12,k_{\Delta}}\approx 0$, and $\Gamma_{1,k_{\Delta}}\Gamma_{2,k_{\Delta}}-\Gamma_{12,k_{\Delta}}^{2}\approx 0$ for $\delta=\pm 0.5$. Therefore, the reflection spectrum exhibits the Lorentzian line shapes centered at $\Delta_{k}=-0.96\Gamma_{e}$ ($0.46\Gamma_{e}$) with the linewidth of $4\Gamma_{e}$, as shown by the red solid curves in Figs. 8(c) and 8(d), respectively. When $\delta=0.5$ ($-0.5$) and $\Delta\approx 1.89J$, the characteristic quantities become $\Delta_{L1,k_{\Delta}}=\Delta_{L2,k_{\Delta}}=\Delta_{L,k_{\Delta}}\approx 0.98\Gamma_{e}$ ($0.83\Gamma_{e}$), $\Gamma_{1,k_{\Delta}}+\Gamma_{2,k_{\Delta}}=4\Gamma_{e}$, $g_{12,k_{\Delta}}\approx 0$, and $\Gamma_{1,k_{\Delta}}\Gamma_{2,k_{\Delta}}-\Gamma_{12,k_{\Delta}}^{2}\approx 0$, which satisfy Eq. (46), as shown by the purple dot-dashed curve in Fig. 8(c) [Fig. 8(d)]. In other parameters, the reflection spectra have two completely reflection peaks, such as $\Delta=1.40J$ (the blue dashed curves). In particular, when $\delta=0.5$ $(-0.5)$ and $\Delta\approx 1.58J$, we have $\Delta_{\text{S}L,k_{\Delta}}\approx\Delta_{\text{A}L,k_{\Delta}}\approx 0$, $\Gamma_{\text{S},k_{\Delta}}\approx 0$, $\Gamma_{\text{A},k_{\Delta}}\approx 7.79\Gamma_{e}$ $(5.26\Gamma_{e})$, $g_{\text{SA},k_{\Delta}}\approx-0.32\Gamma_{e}$ $(-0.95\Gamma_{e})$, and $\Gamma_{\text{SA},k_{\Delta}}\approx 0$. Here, the exchange interaction strength $g_{\text{SA},k_{\Delta}}$ and the individual decay rate of the antisymmetric state $\Gamma_{\text{A},k_{\Delta}}$ satisfy the condition $|g_{\text{SA},k_{\Delta}}|<\Gamma_{\text{A},k_{\Delta}}/4$ Feng2021 ; Abi-Salloum2010 ; Anisimov2011 . In this case, the system exhibits an electromagnetically induced transparency (EIT)-like spectrum. Thus we find that the reflection spectrum has a minimal value ($R_{2}\approx 0$) at $\Delta_{k}\approx 0$ and that the reflection spectrum has two symmetric complete-reflection peaks ($R_{2}=1$) at $\Delta_{k}\approx\pm 0.32\Gamma_{e}$ (±$(\pm 0.95\Gamma_{e})$, as shown by the black dotted curve in Fig. 8(c) [Fig. 8(d)].
Figure 9: Reflection coefficient $R_{2}$ for the AAAB-coupling configuration as a function of $\Delta_{k}$ and $\Delta$. The dimerization parameter is (a) $\delta=0.5$ and (b) $\delta=-0.5$. The profiles of panels (a) and (b) at several values of $\Delta$ are shown by the curves in (c) and (d). Other parameters used are $g=0.01J$ and $d=2$.
@@ III.3.5 The AAAB-coupling configuration
For the AAAB-coupling configuration, the dimensionless parameters in Eqs. (35a)–(36f) are given by $\mu_{1-3}=\nu_{4}=1$ and $\nu_{1-3}=\mu_{4}=0$. In Figs. 9(a) and 9(b), we show the reflection coefficient as a function of $\Delta_{k}$ and $\Delta$. We find that the reflection spectra have two reflection peaks ($R_{2}=1$) at $\Delta_{k}=\Delta_{L1,k_{\Delta}}=\Gamma_{e}\sin(k_{\Delta}d)$ (the first reflection peak) and $\Delta_{k}=\Delta_{L2,k_{\Delta}}=\Gamma_{e}\sin(k_{\Delta}d-\phi_{k_{\Delta}})$ (the second reflection peak). Here, the second Lamb shift $\Delta_{L2,k_{\Delta}}$ depends on $\phi_{k_{\Delta}}$ and hence the position of the second reflection peak is influenced by the sign of $\delta$. To show the details clearly, in Figs. 9(c) and 9(d) we plot the profiles of Figs. 9(a) and 9(b) at different frequency detunings in the region $\Delta\in(J,2J)$, respectively.
When $\delta=\pm 0.5$ and $\Delta\approx 1.58J$ (the blue dashed curves), the reflection spectra exhibit the Lorentzian line shapes. Similarly, we substitute $\delta=\pm 0.5$ and $\Delta\approx 1.58J$ into Eqs. (35b)–(36f) and obtain the reflection coefficient as
@@
Here, we have $\Delta_{L1,k_{\Delta}}=0$, $\Delta_{L2,k_{\Delta}}\approx-0.32\Gamma_{e}$ ($-0.95\Gamma_{e}$), $\Gamma_{1,k_{\Delta}}=0$, $\Gamma_{2,k_{\Delta}}\approx 3.9\Gamma_{e}$ ($2.6\Gamma_{e}$), $g_{12,k_{\Delta}}\approx 0$, and $\Gamma_{12,k_{\Delta}}\approx 0$ for $\delta=0.5$ ($-0.5$). Therefore, the first giant atom is decoupled from the waveguide, while the second giant atom is coupled to the waveguide. In this case, the reflection spectrum exhibits the Lorentzian line shape centered at $\Delta_{k}=-0.32\Gamma_{e}$ ($-0.95\Gamma_{e}$) with the linewidth of $3.9\Gamma_{e}$ ($2.6\Gamma_{e}$), as shown by the blue dashed curve in Fig. 9(c) [Fig. 9(d)]. In particular, when $\delta=\pm 0.5$, $\Delta\approx 1.12J$ (the red solid curves), $1.62J$ (the black dotted curves), and $1.70J$ (the purple dot-dashed curves), the number of the reflection peak depends on the sign of the dimerization parameter $\delta$. Specifically, when $\Delta\approx 1.12J$ (the red solid curves) and $1.70J$ (the purple dot-dashed curves); the reflection spectra at $\delta>0$ have two reflection peaks, while the reflection spectra at $\delta<0$ have one reflection peak and vice versa at $\Delta\approx 1.62J$ (the black dotted curves). This is because the two reflection peaks at $\Delta_{L1,k_{\Delta}}$ and $\Delta_{L2,k_{\Delta}}$ coincide when $\delta=-0.5$ and $\Delta\approx 1.12J$ or $1.70J$ (when $\delta=0.5$ and $\Delta\approx 1.62J$), i.e., $\Delta_{k}=\Gamma_{e}\sin(k_{\Delta}d)=\Gamma_{e}\sin(k_{\Delta}d-\phi_{k_{\Delta}})$. In this case, the reflection spectra have one reflection peak. These analyses indicate that only the position of the second peak depends on the sign of $\delta$. When the sign of $\delta$ changes, the position of the second peak changes accordingly.
Figure 10: Reflection coefficient $R_{2}$ for the AABA-coupling configuration as a function of $\Delta_{k}$ and $\Delta$. The dimerization parameter is (a) $\delta=0.5$ and (b) $\delta=-0.5$. The profiles of panels (a) and (b) at several values of $\Delta$ are shown by the curves in (c) and (d). Other parameters used are $g=0.01J$ and $d=2$.
@@ III.3.6 The AABA-coupling configuration
For the AABA-coupling configuration, the dimensionless parameters in Eqs. (35a)–(36f) are $\mu_{1,2,4}=\nu_{3}=1$ and $\nu_{1,2,4}=\mu_{3}=0$. Figures 10(a) and 10(b) show the reflection coefficient as a function of $\Delta_{k}$ and $\Delta$. It can be found that the reflection spectra have two peaks at $\Delta_{k}=\Gamma_{e}\sin(k_{\Delta}d)$ (the first reflection peak) and $\Delta_{k}=\Gamma_{e}\sin(k_{\Delta}d+\phi_{k_{\Delta}})$ (the second reflection peak), respectively. Therefore, only the position of the second peak depends on the sign of $\delta$, which is similar to the AAAB-coupling configuration. In Figs. 10(c) and 10(d) we show the profiles of Figs. 10(a) and 10(b) at different frequency detunings in the region $\Delta\in(J,2J)$, respectively.
@@
When $\delta=\pm 0.5$ and $\Delta\approx 1.58J$ (the black dotted curves), the reflection spectra exhibit the Lorentzian line shapes, which are consistent with the AAAB-coupling configuration. According to Eqs. (35b)–(36f), we can obtain the reflection coefficient $R_{2}^{\text{AABA}}$, which satisfies Eq. (47). Here, the characteristic quantities become $\Delta_{L1,k_{\Delta}}=0$, $\Delta_{L2,k_{\Delta}}\approx 0.32\Gamma_{e}$ ($0.95\Gamma_{e}$), $\Gamma_{1,k_{\Delta}}=0$, $\Gamma_{2,k_{\Delta}}\approx 3.9\Gamma_{e}$ ($2.6\Gamma_{e}$), $g_{12,k_{\Delta}}\approx 0$, and $\Gamma_{12,k_{\Delta}}\approx 0$ for $\delta=0.5$ ($-0.5$). Therefore, the first giant atom is decoupled from the waveguide, while the second giant atom is coupled to the waveguide. In this case, the reflection spectrum exhibits the Lorentzian line shape centered at $\Delta_{k}=0.32\Gamma_{e}$ ($0.95\Gamma_{e}$) with the linewidth of $3.9\Gamma_{e}$ ($2.6\Gamma_{e}$), as shown by the black dotted curve in Fig. 10(c) [Fig. 10(d)]. From the above analyses, it can be seen that, when $\delta=0.5$ ($-0.5$) and $\Delta\approx 1.58J$, the reflection spectra of AAAB- and AABA-coupling configurations are symmetric with respect to $\Delta_{k}=0$, as shown by the blue dashed curve in Fig. 9(c) [Fig. 9(d)] and the black dotted curve in Fig. 10(c) [Fig. 10(d)]. For other values of $\Delta$, such as $\Delta=1.70J$, the reflection spectra have two reflection peaks [see the purple dot-dashed curves in Figs. 10(c) and 10(d)]. In particular, when $\delta=\pm 0.5$, $\Delta\approx 1.38J$ (the red solid curves), and $1.54J$ (the blue dashed curves), the number of the reflection peaks depends on the sign of $\delta$. Specifically, when $\Delta\approx 1.38J$ (the red solid curves), the reflection spectrum at $\delta>0$ has two reflection peaks, while the reflection spectrum at $\delta<0$ has one reflection peak and vice versa at $\Delta\approx 1.54J$ (the blue dashed curves). When $\delta=-0.5$ and $\Delta\approx 1.38J$ ($\delta=0.5$ and $\Delta\approx 1.54J$), we have the relation $\Delta_{k}=\Gamma_{e}\sin(k_{\Delta}d)=\Gamma_{e}\sin(k_{\Delta}d+\phi_{k_{\Delta}})$. Thus the reflection spectra have one reflection peak, which is similar to the AAAB-coupling configuration.